# m-KAILIN: A Knowledge-Driven Agentic Scientific Corpus Distillation Framework for Biomedical Large Language Models Training

### v2

@@ Abstract
Corpus distillation for biomedical large language models (LLMs) seeks to address the pressing challenge of insufficient quantity and quality in open-source annotated scientific corpora, which remains a bottleneck for effective LLM training in biomedical research. This paper proposes a knowledge-driven, agentic framework for scientific corpus distillation, tailored explicitly for LLM training in the biomedical domain, addressing the challenge posed by the complex hierarchy of biomedical knowledge. Central to our approach is a collaborative multi-agent architecture, where specialized agents—each guided by the Medical Subject Headings (MeSH) hierarchy—work in concert to autonomously extract, synthesize, and self-evaluate high-quality textual data from vast scientific literature. This agentic framework collectively generates and refines domain-specific question-answer pairs, ensuring comprehensive coverage and consistency with biomedical ontologies while minimizing manual involvement. Extensive experimental results show that language models trained on our multi-agent distilled datasets achieve notable improvements in biomedical question-answering tasks, outperforming both strong life sciences LLM baselines and advanced proprietary models. Notably, our AI-Ready dataset enables Llama3-70B to surpass GPT-4 with MedPrompt and Med-PaLM-2, despite their larger scale. Detailed ablation studies and case analyses further validate the effectiveness and synergy of each agent within the framework, highlighting the potential of multi-agent collaboration in biomedical LLM training.
{IEEEkeywords}
Biomedical large language models, corpus distillation, question-answer, agentic AI
@@ 1 Introduction
The advent of large language models has propelled bioinformatics into a new era, enabling the development of automated solutions across a spectrum of biomedical domains [1, 2, 3], and has demonstrated notable success for real-world biomedical question answering (QA) tasks [4, 5]. However, the intricate and specialized nature of biomedical tasks means that general-purpose LLMs often fall short unless meticulously adapted and fine-tuned for the domain [6, 7, 8]. Progress in this area is further constrained by the scarcity of sufficiently large and high-quality biomedical corpora [9, 10, 11]. While existing open-source biomedical datasets are typically of high quality, their limited scale and narrow topical coverage restrict their utility for comprehensive LLM training. Conversely, directly leveraging the expansive body of domain-specific scientific literature offers the potential to cover a broader range of biomedical topics. Nevertheless, most of these resources are unannotated, and their inherent lack of structure, coupled with the absence of QA-format organization, greatly hinders their effective use for training question-answering models. Although the vast body of biomedical literature represents a valuable and authoritative resource, its complex terminology and dense conceptual structures pose significant barriers to automated processing and dataset construction [12, 13]. Figure 1 visualizes the resulting data bottleneck. These challenges raise a crucial question: How can we automatically distill large-scale, high-quality QA pairs from scientific literature to empower biomedical LLM?
@@
Prior research can be categorized into three mainstream categories. First, rule-based methods[14, 15, 16] rely on human-crafted standards for data cleaning and curation, which, while effective in reducing noise, are resource-intensive and difficult to scale. Second, knowledge graph-based approaches[17, 18] structure biomedical information from texts into comprehensive graphs, but their dependence on manually curated sources limits both efficiency and scalability. Third, synthetic data generation methods[19, 20, 21, 22, 23] use LLMs to automate the creation of QA pairs and process large corpora. Yet, they often lack mechanisms for interdisciplinary collaboration[24, 25], resulting in insufficient diversity and robustness in the distilled data. Recent efforts [26] have introduced a knowledge hierarchy-guided [27] approach that leverages a single LLM agent to generate and evaluate biomedical QA data with improved alignment to domain ontologies. Its reliance on a single-agent, rule-based architecture presents inherent limitations in terms of diversity, cross-domain expertise, and collaborative reasoning [28].
@@ 2 Related Work
Dataset distillation[29, 30] is an information extraction technique that uses generative models to distill core information from large-scale raw datasets, resulting in high-quality datasets. Existing dataset distillation approaches are primarily categorized based on their optimization objectives into three types: performance matching, parameter matching, and distribution matching. Performance matching methods [31, 32, 33, 34] aim to optimize dataset distillation by minimizing the loss of a model trained on the distilled dataset when evaluated on the raw dataset. Parameter matching methods[35, 36, 37, 38] focus on the consistency of trainable model parameters when trained on the distilled dataset compared to the raw dataset. Unlike these two training-based objectives, distribution matching methods[39] analyze the distilled dataset itself, using the distributional consistency between the distilled and raw datasets as the target for optimization.
@@
Existing methods for generating and utilizing synthetic text data in biomedical text mining face several challenges. PubMedQA [40] relies on converting titles into questions, often introducing noise due to simplistic or irrelevant content. Attempts to improve PubMedQA through LLM-based rewriting strategies, like those using GPT-3.5-turbo or GPT-4, fail to offer significant diversity in generated content[41]. MedSyn[19], while leveraging a Medical Knowledge Graph and LLMs like GPT-4, lacks a comprehensive integration of hierarchical knowledge. Some other methods that rely on experts to use LLMs with their own knowledge [42, 20] place too much dependence on expert effort, lacking automation and efficiency. Besides, KAILIN [26] first introduced the MeSH knowledge hierarchy as a rule-based evaluator, enhancing the quality of the generated questions. In this study, we advance beyond KAILIN by proposing an agentic, collaborative framework, where multiple specialized agents—each guided by biomedical ontologies—autonomously generate, evaluate, and refine the generated dataset.
@@
Recent advancements [43, 44] in biomedical language models have focused on improving domain adaptation to better handle dense terminology and complex medical concepts. Approaches such as retrieval-augmented generation have enhanced models’ ability to retrieve relevant information for domain-specific tasks[45, 46, 47]. PMC-LLaMA[17] has been fine-tuned using millions of biomedical papers and medical textbooks, integrating medical knowledge graphs like UMLS for domain knowledge. Similarly, BioMistral[14] has applied extensive data processing and adaptive pretraining on PubMed Central, while HEAL[48] has combined public and proprietary datasets for continuous pretraining on a general-purpose model, further enhancing its biomedical capabilities.
@@ 3.1.1 Fine-Tuning on Open-Source QA Dataset
To enable effective biomedical question generation, we fine-tune a large language model on the BioASQ QA dataset [49], and resulting in a specialized generator, $\theta$. Let $\mathcal{T}=\{(d_{i},q_{i})\}_{i=1}^{N}$ denote the open-source QA training set, where $d_{i}$ is a document and $q_{i}$ its reference question. The objective function is:
@@ 3.2 Context Retrieval Agent
The Context Retrieval Agent retrieves relevant supporting documents from the large-scale PubMed [^1] PubMed: Link corpus for each generated question. This agent employs Dense Passage Retrieval (DPR) [50], in line with Retrieval-Augmented Generation (RAG) [51] frameworks. For each question $q_{i}$, we encode both the question and every candidate document $c$ as dense vectors, $\mathbf{e}(q_{i})$ and $\mathbf{e}(c)$, using a shared BiomedBERTbase encoder [^2] BiomedBERT: Link. The relevance score is computed as:
@@ 3.3 Question Evaluation Agent
The Question Evaluation Agent is designed to assess which of two candidate question-context pairs, $(q^{a}_{i},\textbf{c}^{a}_{i})$ and $(q^{b}_{i},\textbf{c}^{b}_{i})$, generated from the same source document $d_{i}$, better aligns with the hierarchical biomedical knowledge embodied in Medical Subject Headings [52] (MeSH). The agent employs a two-stage approach: an initial rule-based evaluation leveraging the MeSH ontology to create high-quality preference labels, followed by learning an automatic evaluator via LLM fine-tuning.
@@ 3.3.1 Knowledge-Driven Fine-Tuning
To enable evaluation without relying on human-annotated preference data, we propose a knowledge-guided, rule-based cold-start method to compare candidate question-context pairs. This approach leverages the structural and semantic information from the MeSH ontology to generate high-confidence preference labels at scale. Given a source document $d_{i}$ and two candidate pairs $(q^{a}_{i},\textbf{c}^{a}_{i})$ and $(q^{b}_{i},\textbf{c}^{b}_{i})$, we can obtain MeSH terms [^3] Our dataset comprises research articles from PubMed, each accompanied by its associated MeSH terms for both the document and its context. from $d_{i}$, $\textbf{c}^{a}_{i}$, and $\textbf{c}^{b}_{i}$. Referring to Lin’s approach [53], for given context set $\textbf{c}_{i}$, the knowledge hierarchy similarity to $d_{i}$ is computed as:
@@ 3.4 Answer Generation Agent
Once the optimal question-context pair is identified, the Answer Generation Agent $\varphi$ generates high-quality answers using advanced LLMs such as GPT-4o [54]. For each input $(q_{j},\textbf{c}_{j})$, this agent produces an answer $\varphi(q_{j},\textbf{c}_{j})\rightarrow a_{j}$, facilitating the construction of comprehensive QA datasets for downstream training.
@@ 3.5 Agentic Collaborative Framework
This stage integrates outputs from the multiple agents to construct datasets, and further optimizes the question generator through direct preference optimization [55] (DPO):
@@ 3.5.1 Initialize Two Distinct Question Generation Agents
As depicted in Figure 2, m-KAILIN framework will first initialize two Question Generation Agents. Specifically, one of them, $\theta^{a}$, will be based on a domain-specific LLM, e.g., BioMistral [14]. And the other $\theta^{b}$ will be based on a more general yet powerful LLM, e.g., Llama-3 [56]. Each of them will follow the same fine-tuning step introduced in Section 3.1.1 and be able to generate quality questions. We want to highlight that the reason we adopt two distinct Question Generation Agents is to enhance diversity and robustness in question generation. The domain-specific agent (e.g., BioMistral) is expected to capture fine-grained biomedical knowledge and terminology, ensuring that generated questions are highly relevant to the biomedical context. Meanwhile, the general-purpose agent (e.g., Llama-3) possesses broader reasoning and linguistic capabilities and may generate questions from more diverse perspectives. By leveraging both agents, m-KAILIN can comprehensively explore the question space over biomedical documents, increasing the likelihood of producing high-quality, informative, and non-redundant question candidates.
@@ 3.6 Training for Downstream Tasks
Finally, the constructed high-quality training corpus is employed to train a target large language model for biomedical question answering in two sequential stages. For instance, we denote the downstream target model as BioMistral [14].
{bclogo}
[…]
@@
{bclogo}
[…]
@@ 3.6.1 Continuous Pre-training
(a) LLMs of fewer than 13B parameters
(b) LLMs of more than 70B parameters
Figure 7: Evaluations (accuracy (%)) for PubMedQA [40] problems on our models compared to other open-source models, closed-source models, and domain-specific models. Models marked with * indicate that they are domain-specific large models focused on the biomedical field rather than general-purpose models. The # symbol denotes closed-source models as opposed to open-source ones.
@@ 3.6.2 Supervised Fine-tuning
Subsequently, the model is fine-tuned on question-context-answer triples $\mathcal{I}_{\textsc{sft}}=\{(q_{j},\textbf{c}_{j},a_{j})\}_{j=1}^{N}$, as well as optionally on benchmark QA datasets such as PubMedQA PQA-A [40] training set. As shown in Figure 6, this prompt is designed to match the downstream QA format. This stage explicitly optimizes the model for accurate answer generation and robust question understanding.
@@ 4.1.1 Base Models
We utilized the llama-series LLM as base models in our primary experiment, while also incorporating BioMistral [14] as the backend of the Question Generation Agent for building the preference dataset and fine-tuning the question generator. We adopted GPT-4o as the backend of the Answer Generation Agent.
@@ 4.1.2 Baselines
We conducted a comprehensive evaluation of various open-source models, including LLaMA-2 [57], LLaMA-3 [56], Mistral [58], GLM-4 [59], Qwen2.5 [60], and Gemma[61], Deepseek [62], as well as proprietary models like GPT-4[54] (with MedPrompt [63, 64]), and Med-PaLM [65]. In particular, we focused on models specifically trained for the biomedical domain, such as BioMistral [14], PMC-LLaMA [17], HEAL [48], KAILIN [26] (with Qwen2.5, GLM-4, and LLama-3 as backend), and MMedLM [66], to demonstrate the effectiveness.
4.1.3 Evaluation Datasets
We validated the results of our main experiment using the PubMedQA benchmark [40], a dataset specifically designed to assess the performance of question-answering systems in the biomedical domain. PubMedQA is tailored to address questions relevant to biomedical literature, making it highly suitable for assessing our framework’s adaptation in this field. Additionally, we categorized the benchmark based on Medical Subject Headings (MeSH) [52] and publication dates, enabling us to evaluate our system’s improved understanding and robustness across diverse biomedical topics and varying time spans:
  • Chronological Subsets for Robustness Evaluation: We evaluated chronological robustness by dividing the PubMedQA PQA-L dataset into 8 time-based subsets, each representing a distinct publication period, to test model performance across different eras.
  • Subdisciplinary Subsets for Robustness Evaluation: We divided the PQA-L dataset into 6 overlapping subsets based on the most common MeSH terms to assess the m-KAILIN framework’s robustness across subdisciplines.

### v3

# Knowledge-Driven Agentic Scientific Corpus Distillation Framework for Biomedical Large Language Models Training

@@ Abstract
Corpus distillation for biomedical large language models (LLMs) seeks to address the pressing challenge of insufficient quantity and quality in open-source annotated scientific corpora, which remains a bottleneck for effective LLM training in biomedical research. This paper proposes a knowledge-driven, agentic framework for scientific corpus distillation, tailored explicitly for LLM training in the biomedical domain, addressing the challenge posed by the complex hierarchy of biomedical knowledge. Central to our approach is a collaborative multi-agent architecture, where specialized agents—each guided by the Medical Subject Headings (MeSH) hierarchy—work in concert to autonomously extract, synthesize, and self-evaluate high-quality textual data from vast scientific literature. This agentic framework collectively generates and refines domain-specific question-answer pairs, ensuring comprehensive coverage and consistency with biomedical ontologies while minimizing manual involvement. Extensive experimental results show that language models trained on our multi-agent distilled datasets achieve notable improvements in biomedical question-answering tasks, outperforming both strong life sciences LLM baselines and advanced proprietary models. Notably, our AI-Ready dataset enables Llama3-70B to surpass GPT-4 with MedPrompt and Med-PaLM-2, despite their larger scale. Detailed ablation studies and case analyses further validate the effectiveness and synergy of each agent within the framework, highlighting the potential of multi-agent collaboration in biomedical LLM training. [^1] The code supporting the findings of this study is available at DropBox.
keywords:
Biomedical Large Language Models, Agentic Corpus Distillation, Synthetic Question–Answer Generation, Agentic AI, Knowledge Hierarchy Guidance
@@ 1 Introduction
The advent of large language models has propelled bioinformatics into a new era [xi2025rise], enabling the development of automated solutions across a spectrum of biomedical domains [luo2025intention, intro-1-1, intro-2-1, 10878496], and has demonstrated notable success for real-world biomedical question answering (QA) tasks [intro-2-2, wang2024biorag, xie2017topic]. However, the intricate and specialized nature of biomedical tasks means that general-purpose LLMs often fall short unless meticulously adapted and fine-tuned for the domain [intro-3-3, icde1, qin2025scihorizon]. Progress in this area is further constrained by the scarcity of sufficiently large and high-quality biomedical corpora [intro-1-2, intro-1-3, ai2025distillingclosedsourcellmsknowledge]. While existing open-source biomedical datasets are typically of high quality, their limited scale and narrow topical coverage restrict their utility for comprehensive LLM training. Conversely, directly leveraging the expansive body of domain-specific scientific literature offers the potential to cover a broader range of biomedical topics. Nevertheless, most of these resources are unannotated, and their inherent lack of structure, coupled with the absence of QA-format organization, greatly hinders their effective use for training question-answering models. Although the vast body of biomedical literature represents a valuable and authoritative resource, its complex terminology and dense conceptual structures pose significant barriers to automated processing and dataset construction [intro-3-2, intro-3-1]. Figure 1 visualizes the resulting data bottleneck. These challenges raise a crucial question: How can we automatically distill large-scale, high-quality QA pairs from scientific literature to empower biomedical LLM?
@@
Prior research can be categorized into three mainstream categories. First, rule-based methods[labrak2024biomistral, data_process-1, yangdistribution] rely on human-crafted standards for data cleaning and curation, which, while effective in reducing noise, are resource-intensive and difficult to scale. Second, knowledge graph-based approaches[wu2024pmc, intro-prior-kg] structure biomedical information from texts into comprehensive graphs, but their dependence on manually curated sources limits both efficiency and scalability. Third, synthetic data generation methods[medsyn-intro-prior-related-syn-2, med-text-mining, cai2023resolving, icde2, xiao2025interdisciplinary] use LLMs to automate the creation of QA pairs and process large corpora. Yet, they often lack mechanisms for interdisciplinary collaboration[10052697, intro-isolate-source-bad], resulting in insufficient diversity and robustness in the distilled data. Recent efforts [cai2025knowledge] have introduced a knowledge hierarchy-guided [xiao2021expert] approach that leverages a single LLM agent to generate and evaluate biomedical QA data with improved alignment to domain ontologies. Its reliance on a single-agent, rule-based architecture presents inherent limitations in terms of diversity, cross-domain expertise, and collaborative reasoning [luo2025large].
@@ 2 Related Work
Dataset distillation[DD, fang2025knowledge] is an information extraction technique that uses generative models to distill core information from large-scale raw datasets, resulting in high-quality datasets. Existing dataset distillation approaches are primarily categorized based on their optimization objectives into three types: performance matching, parameter matching, and distribution matching. Performance matching methods [DD_performance_1, DD_performance_2, DD_performance_3, DD_performance_4] aim to optimize dataset distillation by minimizing the loss of a model trained on the distilled dataset when evaluated on the raw dataset. Parameter matching methods[DD_parameter_1, DD_parameter_2, DD_parameter_3, DD_parameter_4] focus on the consistency of trainable model parameters when trained on the distilled dataset compared to the raw dataset. Unlike these two training-based objectives, distribution matching methods[DD_distribution_1] analyze the distilled dataset itself, using the distributional consistency between the distilled and raw datasets as the target for optimization.
@@
Existing methods for generating and utilizing synthetic text data in biomedical text mining face several challenges. PubMedQA [jin2019pubmedqa] relies on converting titles into questions, often introducing noise due to simplistic or irrelevant content. Attempts to improve PubMedQA through LLM-based rewriting strategies, like those using GPT-3.5-turbo or GPT-4, fail to offer significant diversity in generated content[improve-pubmedqa]. MedSyn[medsyn-intro-prior-related-syn-2], while leveraging a Medical Knowledge Graph and LLMs like GPT-4, lacks a comprehensive integration of hierarchical knowledge. Some other methods that rely on experts to use LLMs with their own knowledge [twelve-tips, med-text-mining] place too much dependence on expert effort, lacking automation and efficiency. Besides, KAILIN [cai2025knowledge] first introduced the MeSH knowledge hierarchy as a rule-based evaluator, enhancing the quality of the generated questions. In this study, we advance beyond KAILIN by proposing an agentic, collaborative framework, where multiple specialized agents—each guided by biomedical ontologies—autonomously generate, evaluate, and refine the generated dataset.
@@
Recent advancements [10614318, 10767279] in biomedical language models have focused on improving domain adaptation to better handle dense terminology and complex medical concepts. Approaches such as retrieval-augmented generation have enhanced models’ ability to retrieve relevant information for domain-specific tasks[intro-4-1, intro-4-2, intro-4-3]. PMC-LLaMA[wu2024pmc] has been fine-tuned using millions of biomedical papers and medical textbooks, integrating medical knowledge graphs like UMLS for domain knowledge. Similarly, BioMistral[labrak2024biomistral] has applied extensive data processing and adaptive pretraining on PubMed Central, while HEAL[HEAL] has combined public and proprietary datasets for continuous pretraining on a general-purpose model, further enhancing its biomedical capabilities.
@@ 3.1.1 Fine-Tuning on Open-Source QA Dataset
To enable effective biomedical question generation, we fine-tune a large language model on the BioASQ QA dataset [bioasq-3], and resulting in a specialized generator, $\theta$. Let $\mathcal{T}=\{(d_{i},q_{i})\}_{i=1}^{N}$ denote the open-source QA training set, where $d_{i}$ is a document and $q_{i}$ its reference question. The objective function is:
@@ 3.2 Context Retrieval Agent
The Context Retrieval Agent retrieves relevant supporting documents from the large-scale PubMed [^2] PubMed: Link corpus for each generated question. This agent employs Dense Passage Retrieval (DPR) [DPR], in line with Retrieval-Augmented Generation (RAG) [RAG] frameworks. For each question $q_{i}$, we encode both the question and every candidate document $c$ as dense vectors, $\mathbf{e}(q_{i})$ and $\mathbf{e}(c)$, using a shared BiomedBERTbase encoder [^3] BiomedBERT: Link. The relevance score is computed as:
@@ 3.3 Question Evaluation Agent
The Question Evaluation Agent is designed to assess which of two candidate question-context pairs, $(q^{a}_{i},\textbf{c}^{a}_{i})$ and $(q^{b}_{i},\textbf{c}^{b}_{i})$, generated from the same source document $d_{i}$, better aligns with the hierarchical biomedical knowledge embodied in Medical Subject Headings [mesh] (MeSH). The agent employs a two-stage approach: an initial rule-based evaluation leveraging the MeSH ontology to create high-quality preference labels, followed by learning an automatic evaluator via LLM fine-tuning.
@@ 3.3.1 Knowledge-Driven Fine-Tuning
To enable evaluation without relying on human-annotated preference data, we propose a knowledge-guided, rule-based cold-start method to compare candidate question-context pairs. This approach leverages the structural and semantic information from the MeSH ontology to generate high-confidence preference labels at scale. Given a source document $d_{i}$ and two candidate pairs $(q^{a}_{i},\textbf{c}^{a}_{i})$ and $(q^{b}_{i},\textbf{c}^{b}_{i})$, we can obtain MeSH terms [^4] Our dataset comprises research articles from PubMed, each accompanied by its associated MeSH terms for both the document and its context. from $d_{i}$, $\textbf{c}^{a}_{i}$, and $\textbf{c}^{b}_{i}$. Referring to Lin’s approach [sim_lin], for given context set $\textbf{c}_{i}$, the knowledge hierarchy similarity to $d_{i}$ is computed as:
@@ 3.4 Answer Generation Agent
Once the optimal question-context pair is identified, the Answer Generation Agent $\varphi$ generates high-quality answers using advanced LLMs such as GPT-4o [achiam2023gpt]. For each input $(q_{j},\textbf{c}_{j})$, this agent produces an answer $\varphi(q_{j},\textbf{c}_{j})\rightarrow a_{j}$, facilitating the construction of comprehensive QA datasets for downstream training.
@@ 3.5 Agentic Collaborative Framework
This stage integrates outputs from the multiple agents to construct datasets, and further optimizes the question generator through direct preference optimization [DPO] (DPO):
@@ 3.5.1 Initialize Two Distinct Question Generation Agents
As depicted in Figure 2, m-KAILIN framework will first initialize two Question Generation Agents. Specifically, one of them, $\theta^{a}$, will be based on a domain-specific LLM, e.g., BioMistral [labrak2024biomistral]. And the other $\theta^{b}$ will be based on a more general yet powerful LLM, e.g., Llama-3 [llama3_1]. Each of them will follow the same fine-tuning step introduced in Section 3.1.1 and be able to generate quality questions. We want to highlight that the reason we adopt two distinct Question Generation Agents is to enhance diversity and robustness in question generation. The domain-specific agent (e.g., BioMistral) is expected to capture fine-grained biomedical knowledge and terminology, ensuring that generated questions are highly relevant to the biomedical context. Meanwhile, the general-purpose agent (e.g., Llama-3) possesses broader reasoning and linguistic capabilities and may generate questions from more diverse perspectives. By leveraging both agents, m-KAILIN can comprehensively explore the question space over biomedical documents, increasing the likelihood of producing high-quality, informative, and non-redundant question candidates.
@@ 3.6 Training for Downstream Tasks
Finally, the constructed high-quality training corpus is employed to train a target large language model for biomedical question answering in two sequential stages. For instance, we denote the downstream target model as BioMistral [labrak2024biomistral].
To address the challenges for the biomedical field, I reviewed the paper titled: {Title}, {Context}.
@@
Please analyze the information in the title and context in the field of biomedical and generate a question:
@@ 3.6.2 Supervised Fine-tuning
Subsequently, the model is fine-tuned on question-context-answer triples $\mathcal{I}_{\textsc{sft}}=\{(q_{j},\textbf{c}_{j},a_{j})\}_{j=1}^{N}$, as well as optionally on benchmark QA datasets such as PubMedQA PQA-A [jin2019pubmedqa] training set. As shown in Figure 6, this prompt is designed to match the downstream QA format. This stage explicitly optimizes the model for accurate answer generation and robust question understanding.
@@ 4.1.1 Base Models
We utilized the llama-series LLM as base models in our primary experiment, while also incorporating BioMistral [labrak2024biomistral] as the backend of the Question Generation Agent for building the preference dataset and fine-tuning the question generator. We adopted GPT-4o as the backend of the Answer Generation Agent.
@@ 4.1.2 Baselines
We conducted a comprehensive evaluation of various open-source models, including LLaMA-2 [touvron2023llama], LLaMA-3 [llama3_1], Mistral [jiang2023mistral], GLM-4 [glm2024chatglm], Qwen2.5 [yang2024qwen2], and Gemma[team2024gemma], Deepseek [bi2024deepseek], as well as proprietary models like GPT-4[achiam2023gpt] (with MedPrompt [nori2023can, nori2023capabilities]), and Med-PaLM [Med-Palm-2]. In particular, we focused on models specifically trained for the biomedical domain, such as BioMistral [labrak2024biomistral], PMC-LLaMA [wu2024pmc], HEAL [HEAL], KAILIN [cai2025knowledge] (with Qwen2.5, GLM-4, and LLama-3 as backend), and MMedLM [MMed_LM], to demonstrate the effectiveness.
4.1.3 Evaluation Benchmarks and Protocols
We conduct a comprehensive evaluation using a suite of benchmark subsets derived from PubMedQA [jin2019pubmedqa]. We design a multi-dimensional evaluation protocol that probes complementary aspects of biomedical QA capability, including overall accuracy, temporal generalization, subdisciplinary robustness, reasoning sensitivity, and data scaling behavior. Specifically, we first evaluate all models on the PubMedQA-Full benchmark, which consists of expert-annotated question–answer pairs grounded in biomedical literature and serves as the primary indicator of end-to-end QA performance. To assess robustness under biomedical knowledge evolution, we further construct chronological benchmarks (PubMedQA-Chrono) by partitioning the PQA-L subset into eight time-based slices according to publication periods, enabling evaluation across distinct biomedical eras.
To examine cross-domain generalization, we derive subdisciplinary benchmarks (PubMedQA-MeSH) by grouping samples according to their dominant Medical Subject Headings [mesh] (MeSH) terms, resulting in six overlapping subsets that reflect major biomedical research areas. In addition, following prior work, we distinguish between reasoning-required and question-only inference settings, forming a reasoning-oriented benchmark that isolates the contribution of contextual understanding and biomedical reasoning.
Together, these benchmark settings provide a controlled yet comprehensive evaluation suite, allowing us to systematically analyze the effectiveness and robustness of the proposed m-KAILIN framework across multiple dimensions of biomedical question answering.